Consider a company with the following cost and revenue information: ATC = $80, AVC = $70, and MR = MC = $90. If the firm produces Q = 600 in the short run, it:

Group of answer choices

A. is minimizing losses.

B. makes a total loss of $6000.

C. should produce more output.

D. is making a mistake and should shut down.

E. is maximizing total profit.
Answer #1

Answer

Option E

E. is maximizing total profit.

the firm is maximizing profit as the MR=MC is means the profit is maximum

also, the MR>ATC means the profit is maximum and not loss is minimum because the firm is making profit

Even, we don't know the firm is in perfect competition or imperfect competition but the P equal to or higher than the MR

so the firm is making a minimum profit of

Profit =(P-ATC)*Q

=(90-80)*600=$6000

the firm is making a minimum profit of $6000